if x represents the first, or the smaller, of the two consecutive even integers, express the sum of the two integers in terms of x
Answer by Shana-D77(132)

x is the first integer.
The next consecutive integer is x + 1
The sum of the two integers is:
x + (x + 1)
Simplified, this is:
2x + 1


Let's see if it works with the two consecutive integers 5 and 6:
let x = 5
so x + 1 = 6
5 + 6 = 11, so:
2x + 1 = 11
2x = 10 (subtracted 1 from both sides)
x = 5 (divided both sides by 2)
it works!